OpenStudy (anonymous):

A 100 kg crate is dragged across the floor at a constant velocity by a rope. If the applied force from the rope is 800 N and the force from friction is 500 N, find the angle at which the crate is being pulled.

OpenStudy (akashdeepdeb):

 800 cos (a) - 500 = 0 [Newton's 2nd law of motion] cos (a) = 5/8 a = \(cos^{-1} (\frac{5}{8})\) Understood? :)